Activity-Based Product Costing

Suppose that a surgical ward has gathered the following information for four nursing activities and two types of patients:

Patient Category
Driver Normal Intensive Activity Rate
Treating patients Treatments 7,400 9,500 $4.00
Providing hygienic care Hygienic hours 5,800 18,100 5.00
Responding to requests Requests 32,000 80,000 2.00
Monitoring patients Monitoring hours 6,000 72,000 3.00

Required:

1. Determine the total nursing costs assigned to each patient category.

Costs Assigned
Normal $
Intensive $

2. Output is measured in patient days. Assuming that the normal patient category uses 8,000 patient days and the intensive patient category uses 6,400 patient days, calculate the nursing cost per patient day for each type of patient. Round your answers to the nearest cent.

Normal $ per patient day
Intensive $ per patient day

3. Conceptual Connection: The supervisor of the surgical ward has suggested that patient days is the only driver needed to assign nursing costs to each type of patient. Calculate the charge per patient day using this approach. Round your answer to the nearest cent.
$ per patient day
Answer #1

Driver

Rate

Normal

Intensive

Total

Treating patient

Treatment

4

$    29,600.00

$    38,000.00

$    67,600.00

Providing hygine care

hygiene hours

5

$    29,000.00

$    90,500.00

$ 119,500.00

Responding to request

Requests

2

$    64,000.00

$ 160,000.00

$ 224,000.00

Monitoring Patient

Monitoring Hours

3

$    18,000.00

$ 216,000.00

$ 234,000.00

Total Cost

$ 140,600.00

$ 504,500.00

$ 645,100.00

Ans:

1. Determine the total nursing costs assigned to each patient category.

Calculation of Total Nursing costs Assigned to Each patient Category:

Costs Assigned

Normal

$140,600

Intensive

$504,500

2. Output is measured in patient days. Assuming that the normal patient category uses 8,000 patient days and the intensive patient category uses 6,400 patient days, calculate the nursing cost per patient day for each type of patient. Round your answers to the nearest cent.

Ans:

Normal=$140,600/8000=17.575

$ 17.58 per patient day

Intensive=504,500/6400=78.828

$ 78.83 per patient day

3. Conceptual Connection: The supervisor of the surgical ward has suggested that patient days are the only driver needed to assign nursing costs to each type of patient. Calculate the charge per patient day using this approach. Round your answer to the nearest cent.
$ per patient day

Ans:

Total Patient Days (8000+6400)

               14400

Total Cost

$           645,100.00

Cost per patient days = 645,100/14400

$                      44.80